### v2

# Symmetry-enforced Moiré Topology

Abstract
Topological flat bands in two-dimensional (2D) moiré materials have emerged as promising platforms for exploring the interplay between topology and correlation effects. However, realistic calculations of moiré band topology using density functional theory (DFT) are computationally inefficient due to the large number of atoms in a single moiré unit cell. In this work, we propose a systematic scheme to predict the topology of moiré bands from atomic symmetry data and moiré symmetry group, both of which can be efficiently extracted from DFT. Specifically, for $\Gamma$-valley electron gases, we find that certain combinations of atomic symmetry data and moiré symmetry groups can enforce nontrivial band topology in the low-energy moiré bands, as long as the moiré band gap is smaller than the atomic band splitting at the moiré Brillouin zone boundary. This symmetry-enforced nontrivial moiré topology, including both topological insulators and topological semimetals, is robust against various material-specific details such as the precise form and strength of the moiré potential or the exact twist angle. By exhaustively scanning all 2D atomic symmetry data and moiré symmetry groups, we identify 197 combinations that can yield symmetry-enforced nontrivial moiré topology, and we verify one such combination using a moiré model with cubic Rashba spin-orbit coupling. By screening the existing 2D material database, we currently identify 92 monolayer materials with (i) the low-energy bands near $\Gamma$ and (ii) the atomic symmetry data that belong to those combinations. Our approach is generalizable to other valleys and provides a useful guideline for experimental efforts to discover and design new topologically nontrivial moiré materials.
I Introduction
Moiré materials [cao2018unconventional, bistritzer2011moire] have emerged as one of the most important platforms for exploring exotic strongly correlated physics, primarily because they host energy bands that are (nearly) flat and topological. The most recent prominent example is twisted bilayer MoTe2, which has been extensively studied both experimentally [cai2023signatures, zeng2023thermodynamic, park2023observation, xu2023observation, ji2024local, redekop2024direct, kang2024evidence, park2025ferromagnetism, an2025observation, jia2025anomalous, xu2025interplay, kang2025time, park2025observation, xu2025signatures, xia2025simulating] and theoretically [wu2019topological, yu2020giant, pan2020band, zhang2021electronic, li2021spontaneous, devakul2021magic, morales2023pressure, wang2024fractional, reddy2023fractional, qiu2023interaction, dong2023composite, wang2023diverse, goldman2023zero, morales2024magic, liu2024gate, xu2024maximally, reddy2023toward, song2024phase, wu2024time, yu2024fractional, abouelkomsan2024band, li2024electrically, jia2024moire, mao2024transfer, zhang2024polarization, wang2023topology, li2024contrasting, sheng2024quantum, reddy2024non, xu2025multiple, ahn2024non, wang2025higher, shen2024stabilizing, wang2024phase, kwan2024could, wu2024quantum, zaklama2025structure, zhang2024universal, qiu2025topological, gonccalves2025spinless, liu2025characterization], where the flat bands with nonzero Chern numbers give rise to fractional Chern insulators [neupert2011fractional, sheng2011fractional, regnault2011fractional, tang2011high, sun2011nearly] at fractional fillings. The small bandwidth of the moiré bands in moiré materials naturally arises from the large moiré unit cell, which folds the non-moiré bands (i.e., atomic bands) hundreds of times, leading to energy bands that have bandwidths much smaller than their non-moiré counterparts. However, the nontrivial topology of the flat bands is not necessarily guaranteed. There is one known example of enforcing the nontrivial topology of the low-energy moiré bands based on the non-moiré band topology—twisted bilayer graphene (and related graphene-based systems) [cao2018unconventional, cao2018correlated, bistritzer2011moire, song2019all, po2019faithful, ahn2019failure, tarnopolsky2019origin, song2021twisted, song2022magic]. Yet, this case is quite special: one needs both the $C_{2}\mathcal{T}$ symmetry (i.e., combination of two-fold rotation $C_{2}$ and time reversal (TR) symmetry $\mathcal{T}$) and an effective normal-state particle-hole symmetry [song2021twisted], which, especially the latter, are not common beyond graphene-based systems. Indeed, the nontrivial topology of the flat bands in twisted bilayer MoTe2 is not guaranteed—the Chern numbers of the bands are sensitive to material parameters such as twisted angles [zhang2024polarization, xu2025multiple, zhang2024universal, li2024electrically].
It would be transformative to develop general principles that can guarantee the nontrivial topology of moiré flat bands based on ubiquitous symmetries, such as crystalline and TR symmetries. The reason is summarized in Fig. 1 and elaborated in the following. Given a realistic moiré system formed by 2D non-moiré layered materials, it is extremely time consuming to directly calculate the moiré bands and their band topology (the red box in Fig. 1), especially when the moiré unit cell becomes very large, e.g. twist bilayer MoTe2 with a twist angle smaller than $2^{\circ}$. On the other hand, it is highly efficient to know the symmetry group (crystalline and TR) and symmetry representations of the non-moiré layered materials at high-symmetry momenta (which together we refer to as atomic symmetry data), and it takes no time to know the symmetry group of the moiré system (which we refer to as moiré symmetry group), as depicted by the green boxes in Fig. 1. Therefore, the principles that can derive moiré topology from atomic symmetry data and moiré symmetry group can dramatically advance our ability of numerically predicting new topological moiré platforms. They are also crucial for experimental studies, as those principles should be robust against sizable variations in material parameters that are often difficult to control in practice—such as the twist angle and moiré potential/tunneling strength.
In this work, we successfully establish such a principle. Specifically, for $\Gamma$-valley electron gas subject to weak moiré potential, we find that certain combinations of atomic symmetry group, atomic symmetry representation at $\Gamma$ and moiré symmetry group can enforce nontrivial topology of the low-energy moiré bands, where nontrivial topology includes both topological insulator and topological semi-metal phases. Here the low-energy moiré bands refer to the moiré bands that are closest to the charge neutrality, i.e., conduction bottom bands or valence top bands. We refer to such enforced nontrivial topology as symmetry-enforced moiré topology. Weak moiré potential means that the strength of the moiré potential is weaker than the atomic band splitting, which is a realistic condition at $\Gamma$ valley satisfied by known examples such as twisted bilayer MoTe2, as discussed in Sec.VI. In other words, for those combinations of atomic symmetry data and moiré symmetry group, the phase diagram of the low-energy bands of the moiré model only includes topological insulator and semi-metal phases—no topologically trivial phases. By systematically analyzing all possible combinations, we identify 197 combinations that can lead to symmetry-enforced nontrivial moiré topology, with 16 of them for the topological insulating phase in the phase diagram. The general principle is explicitly verified in a moiré model with cubic Rashba spin-orbit coupling (SOC). We further screen the existing non-moiré 2D material databases [vergniory2019complete, haastrup2018computational, jiang20242d], focusing on monolayers with plane groups P31m, P3m1, and P3. We identify 92 monolayer materials that have (i) low-energy bands near $\Gamma$ and (ii) the atomic symmetry data that belong to those combinations, as discussed in Sec. V.
Clearly, our principle does not rely on precisely-tuned values of moiré parameters such as twist angle and moiré potential strength. It is also independent of the form of moiré potential, in contrast to Refs.[crepel2025efficient, Lhachemiarxiv]. We note that the moiré topological semi-metal phases may also be a fertile ground for exotic strongly-correlated phases. It is because electron-electron interactions may open a gap between topological metallic bands and drive the system into topologically ordered states. A closely related example is the emergence of fractional Chern insulators in multilayer rhombohedral graphene/hexagonal boron nitride (hBN) heterostructures at various electron fillings [lu2024fractional, xie2025tunable, waters2025chern, lu2025extended, aronson2025displacement, zhou2024layer, han2024engineering, xiang2025continuously, wang2025electrical, li2025tunable, xie2025unconventional], where the moiré bands are nearly gapless at the single-particle level [park2023topological, dong2024theory, zhou2024fractional, dong2024anomalous, herzog2024moire, guo2024fractional, kwan2025moire, soejima2024anomalous, dong2024stability, xie2024integer, kudo2024quantum, crepel2025efficient, huang2024self, yu2024moir, huang2025fractional, das2024thermal, wei2025edge, zhou2025new, zeng2025berry, bernevig2025berry, li2025multiband, uchida2025non].
Figure 1: Theoretical Approaches to Identify Moiré Topology. All green blocks are easily computationally accessible, while the red block is computationally inefficient. Our work provides an efficient way to indicate moiré topology from the moiré symmetry group and atomic symemtry data.
Figure 2: Schematics of Symmetry-Enforced Moiré Topology. (a) The band structure around $\Gamma$ point of a pristine layered material with point group $C_{6v}$. The low-energy band at $\Gamma$ is described by the 2D $\bar{\Gamma}_{7}$ irrep. (b) Band folding at zero moiré potential. At ${\text{K}}_{M}$, the folded bands form a three-dimensional representation $\bar{\text{K}}_{4}\oplus\bar{\text{K}}_{6}$, while at ${\text{M}}_{M}$ they form a two-dimensional irrep $\bar{\text{M}}_{5}$. $\Delta_{A}$ marks the energy scale of the atomic band splitting at the mBZ boundary. (c) One possible moiré band structure in the presence of weak moiré potential, i.e., the energy splitting generated by $\Delta_{M}$ is smaller than $\Delta_{A}$. Here the top two bands (red) carry the $\bar{\text{K}}_{6}$ irrep, while the third top band has the $\bar{\text{K}}_{4}$ irrep, resulting in nontrivial topology in the isolated top two bands. (d) The other possible moiré band structure in the presence of weak moiré potential. The $\bar{\text{K}}_{4}$ band has higher energy than the two $\bar{\text{K}}_{6}$ bands, resulting in a topological semi-metallic phase for the red bands due to the unavoidable touching between the second and third top bands at $\text{K}_{M}$.
II Illustration of Principle
We will first use a specific set of atomic symmetry data and moiré symmetry group to illustrate the underlying principle. Let us consider spin-orbit-coupled electrons around $\Gamma$ point of a non-moiré layered material. Suppose the electrons carry total out-of-plane angular momentum $J_{z}=\pm 3/2$, and have the point group $C_{6v}$ and TR symmetry. $J_{z}=\pm 3/2$ can be generated by the addition of atomic orbitals ($p$, $d$ or higher) and electron spin. The $C_{6v}$ group is generated by a six-fold rotation along the out-of-plane direction and a mirror reflection about the $y$-axis ($M_{y}$). Owing to the angular momentum, the electrons at $\Gamma$ form a doublet, furnishing $\bar{\Gamma}_{7}$ irreducible representation (irrep) of $C_{6v}$. Without loss of generality, we choose both bands to bend downward, and the two bands naturally get split by the SOC away from $\Gamma$, as shown in Fig. 2(a).
Now we introduce the moiré potential that preserves the point group $C_{6v}$ and TR symmetry. In addition, the moiré potential would introduce the discrete moiré translation symmetry, leading to a plane group $P6mm$. In order to understand the generation of the moiré bands, let us first consider an artificial limit where the moiré potential strength $\Delta_{M}$ vanishes. In this case, the moiré bands simply come from the artificial band folding of the $\Gamma$-valley atomic bands (Fig. 2(a)) into the moiré Brillouin zone (mBZ), as shown in Fig. 2(b). The $\bar{\Gamma}_{7}$ irrep survives under the band folding and stays at the moiré $\Gamma_{M}$ point, while the band folding generates a Kramers’ pair among the top two bands at the $\text{M}_{M}$ point, furnishing $\bar{\text{M}}_{5}$ irrep. Most importantly, the band folding gives rise to three accidentally degenerate states at $\text{K}_{M}$ point. The degeneracy is accidental because they can be split into a one-dimensional (1D) irrep $\bar{\text{K}}_{4}$ and a two-dimensional (2D) irrep $\bar{\text{K}}_{6}$ according to the moiré symmetry group discussed below.
Now we resume the nonzero moiré strength $\Delta_{M}$. We consider the weak moiré region where the energy splitting induced by the moiré potential is smaller than the atomic band splitting ($\Delta_{A}$) at the mBZ boundary. The moiré potential will leave the $\bar{\Gamma}_{7}$ irrep at $\Gamma_{M}$ and the $\bar{\text{M}}_{5}$ irrep at $\text{M}_{M}$ intact, as protected by symmetry. Nevertheless, the moiré potential will split the $\bar{\text{K}}_{4}$ and $\bar{\text{K}}_{6}$ bands at $\text{K}_{M}$. Focusing on the top two bands, there are two possibilities. One is that the top two bands carry the 2D $\bar{\text{K}}_{6}$ irrep, and the third band possesses the $\bar{\text{K}}_{4}$ irrep, resulting in the top two bands being isolated (Fig. 2(c)). In this case, the top two bands have nontrivial topology as the irreps $\bar{\Gamma}_{7}$, $\bar{\text{M}}_{5}$ and $\bar{\text{K}}_{6}$ can never appear simultaneously for an isolated set of trivial bands according to topological quantum chemistry [bradlyn2017topological, cano2021band, elcoro2021magnetic] and symmetry indicator theory [po2017symmetry, kruthoff2017topological, fu2007topological, fang2012bulk, po2020symmetry, lenggenhager2022universal]. The other possibility is that the top two bands are from one 1D $\bar{\text{K}}_{4}$ irrep and one component of the 2D $\bar{\text{K}}_{6}$ irrep. In this case, the top second band definitely touches the third band as they together form the $\bar{\text{K}}_{6}$ irrep at $\text{K}_{M}$. Therefore, regardless of which possibility, the low-energy moiré bands are either topologically insulating or topologically semi-metallic. Such nontrivial moiré topology is enforced by the atomic symmetry data and moiré symmetry group, as long as the moiré potential is weak compared to atomic band splitting.
III Moiré cubic Rashba model
We now use an specific example to verify the principle. We consider a model Hamiltonian describing a spin-orbit-coupled 2D electron gas around $\Gamma$ point subject to a moiré superlattice potential. As discussed in Sec. II, we impose the point group $C_{6v}$ and TR symmetry, and consider $J_{z}=\pm 3/2$ doublet at $\Gamma$. Then, the atomic part of the Hamiltonian up to the fourth order of the momentum $\mathbf{k}$ has the following general form
$\hat{H}^{A}=\sum_{\mathbf{k}nm}c_{n\mathbf{k}}^{\dagger}[H^{A}(\mathbf{k})]_{nm}c_{m\mathbf{k}},$
$H^{A}(\mathbf{k})=\begin{pmatrix}\alpha k^{2}+\beta k^{4}&iR_{3}k^{3}_{-}\\ -iR_{3}k^{3}_{+}&\alpha k^{2}+\beta k^{4}\end{pmatrix}$. (1)
where $c_{n\mathbf{k}}$ is the fermion annihilation operator with $n=1,2$ indexing $J_{z}=\pm 3/2$. To ensure that both bands of $\hat{H}^{A}$ bend down, we choose both $\alpha$ and $\beta$ to be negative in the kinetic term $\alpha k^{2}+\beta k^{4}$, where the $k^{4}$ term is necessary due to the cubic Rashba SOC term, i.e., the $R_{3}$ term in Sec. III.
We consider the moiré superlattice potential on a hexagonal moiré lattice, which reads
$\hat{H}_{M}=\sum_{\alpha=1,2}\int d^{2}rc_{n\mathbf{r}}^{\dagger}H_{\text{M}}(\mathbf{r})c_{n\mathbf{r}},$ (2)
with $c_{n\mathbf{r}}=\frac{1}{\sqrt{V}}\int_{\mathds{R}^{2}}d^{2}kc_{n\mathbf{k}}e^{i\mathbf{k}\cdot\mathbf{r}}$ and $V$ the system area. We include both the first and second harmonic moiré supperlattice potential, namely, $H_{M}(\mathbf{r})=\sum_{{\boldsymbol{g}}\in\mathbf{G}^{\mathds{1}}_{M}}\Delta_{1}e^{i{\boldsymbol{g}}\cdot\mathbf{r}}+\sum_{{\boldsymbol{g}}\in\mathbf{G}^{\mathbb{2}}_{M}}\Delta_{2}e^{i{\boldsymbol{g}}\cdot\mathbf{r}}$, where $\mathbf{G}^{\mathds{1}}_{M}$ consists of $\mathbf{b}^{M}_{1}$ and its partners under six-fold rotation symmetries, $\mathbf{G}^{\mathbb{2}}_{M}$ consists of $\mathbf{b}^{M}_{1}+\mathbf{b}^{M}_{2}$ and its partners under six-fold rotation symmetries. $\mathbf{b}^{M}_{1}=2\pi(1,-1/\sqrt{3})/a_{M}$ and $\mathbf{b}^{M}_{2}=2\pi(1,1/\sqrt{3})/a_{M}$ are primitive reciprocal lattice vectors, and $a_{M}$ is the moiré lattice constant. The moiré potential in Eq. 2 has plane group $P6mm$, generated by the point group $C_{6v}$ and the moiré lattice translation symmetry, and also preserves the TR symmetry. Those symmetries are preserved even after we incorporate the atomic Hamiltonian in Sec. III, resulting in the total Hamiltonian
$\hat{H}=\hat{H}^{A}+\hat{H}_{M}\ .$ (3)
To illustrate our principle, we focus on the weak moiré region, where the moiré gap is smaller than the band splitting caused by Rashba SOC. By choosing reasonable parameter values (listed in supplementary material (SM) Sec.C1 [SM].), we obtain the phase diagram as a function of $\Delta_{1}$ and $\Delta_{2}$ in Fig. 3(a). The color represents the direct band gap $E_{\Delta}$ between the second and third topmost moiré bands (see Fig. 3(b)) and the red dashed line separates the insulating region from the semi-metallic region for the top two moiré bands. (See more details in SM Sec.C1 [SM].) We find that regardless of the value of the moiré parameter values, there are only two phases for the low-energy physics: (i) topological insulator phase where the top two bands are isolated with band irreps ($\bar{\mathbf{\Gamma}}_{7}$, $\bar{\textbf{M}}_{5}$, $\bar{\textbf{K}}_{6}$) and have $\mathds{Z}_{2}=1$ in Fig. 3(b), and (ii) topological semi-metal phase where the top two bands are enforced to be connected to other bands in Fig. 3(c) as the irrep $\bar{\textbf{K}}_{4}$ for the topmost band is 1D, while $\bar{\textbf{M}}_{5}$ and $\bar{\mathbf{\Gamma}}_{7}$ irreps are 2D. This is consistent with the principle that we discussed in Sec. II.
Figure 3: Cubic Rashba SOC model under moiré potential. (a) Phase diagram of the topmost two bands of the moiré cubic Rashaba model in Eq. 3. $E_{\Delta}$ denotes the band gap between the second and third topmost bands. The red dashed lines separate the insulator region and semimetal region. (b) The band dispersion and band irreps computed with the parameters labeled by the blue cross point in (a). The top two bands have $\mathbb{Z}_{2}=1$. (c) The band dispersion and band irreps computed with the parameters labeled by the red cross point in (a). The top two bands are connected to the top third band.
$G_{0}$ $G^{A}_{0}$ $C_{6v}$ $C_{3v}$ $C_{6}$ $C_{3}$ $C_{4v}$ $C_{2v}$ $C_{s}$
$P6mm$ $\bar{\Gamma}_{7}$
$P3m1$ $\bar{\Gamma}_{7}$ $\bar{\Gamma}_{4}\bar{\Gamma}_{5}$
$P31m$ $\bar{\Gamma}_{7}$ $\bar{\Gamma}_{4}\bar{\Gamma}_{5}$
$P6$ $\bar{\Gamma}_{7}$ $\bar{\Gamma}_{7}\bar{\Gamma}_{8}$
$P3$ $\bar{\Gamma}_{7}$ $\bar{\Gamma}_{4}\bar{\Gamma}_{5}$ $\bar{\Gamma}_{7}\bar{\Gamma}_{8}$ $\bar{\Gamma}_{4}\bar{\Gamma}_{4}$
$P4bm$ $\bar{\Gamma}_{6}$, $\bar{\Gamma}_{7}$
$Pba2$ $\bar{\Gamma}_{7}$, $\bar{\Gamma}_{8}$, $\bar{\Gamma}_{9}$ $\bar{\Gamma}_{6}$, $\bar{\Gamma}_{7}$ $\bar{\Gamma}_{5}$
$Pma2$ $\bar{\Gamma}_{7}$, $\bar{\Gamma}_{8}$, $\bar{\Gamma}_{9}$ $\bar{\Gamma}_{6}$, $\bar{\Gamma}_{7}$ $\bar{\Gamma}_{5}$
$Pb11$ $\bar{\Gamma}_{7}$, $\bar{\Gamma}_{8}$, $\bar{\Gamma}_{9}$ $\bar{\Gamma}_{4}\bar{\Gamma}_{5}$, $\bar{\Gamma}_{6}$ $\bar{\Gamma}_{6}$, $\bar{\Gamma}_{7}$ $\bar{\Gamma}_{5}$ $\bar{\Gamma}_{3}\bar{\Gamma}_{4}$
Table 1: Spinful TR-invariant Cases for Symmetry-enforced Moiré Topology. List of combinations of atomic point group $G_{0}^{A}$, atomic $\Gamma$ irrep, and moiré plane gorup $G_{0}$ that lead to symmetry-enforced nontrivial moiré topology, in the presence of both TR symmetry and SOC. For a given $G_{0}^{A}$ and $G_{0}$, the table entry indicates the required atomic $\Gamma$ irrep to give symmetry-enforced topology. The irrep is highlighted by the red means the low-energy bands can be topological insulating.
$G_{0}$ $G^{A}_{0}$ $C_{6v}$ $C_{4v}$ $C_{3v}$ $C_{2v}$ $C_{s}$
$P6$ ${\Gamma}_{5}$, ${\Gamma}_{6}$
$P2$ ${\Gamma}_{5}$, ${\Gamma}_{6}$ ${\Gamma}_{5}$
$P6mm$ ${{\Gamma}}_{5}$, ${{\Gamma}}_{6}$
$P4mm$ ${{\Gamma}}_{5}$
$P4bm$ ${{\Gamma}}_{1}$, ${{\Gamma}}_{2},{{\Gamma}}_{3}$, ${{\Gamma}}_{4}$
$P3m1$ ${{\Gamma}}_{5}$, ${{\Gamma}}_{6}$
$Pmm2$ ${{\Gamma}}_{5}$, ${{\Gamma}}_{6}$ ${\Gamma}_{5}$
$Cmm2$ ${{\Gamma}}_{5}$, ${{\Gamma}}_{6}$ ${\Gamma}_{5}$
$Pba2$ ${{\Gamma}}_{1}$, ${{\Gamma}}_{2},{{\Gamma}}_{3}$, ${{\Gamma}}_{4}$ ${{\Gamma}}_{1}$, ${{\Gamma}}_{2},{{\Gamma}}_{3}$, ${{\Gamma}}_{4}$ ${{\Gamma}}_{1}$, ${{\Gamma}}_{2},{{\Gamma}}_{3}$, ${{\Gamma}}_{4}$
$Pma2$ ${{\Gamma}}_{1}$, ${{\Gamma}}_{2},{{\Gamma}}_{3}$, ${{\Gamma}}_{4},{{\Gamma}}_{5}$, ${{\Gamma}}_{6}$ ${{\Gamma}}_{1}$, ${{\Gamma}}_{2},{{\Gamma}}_{3}$, ${{\Gamma}}_{4}$, ${{\Gamma}}_{5}$ ${{\Gamma}}_{1}$, ${{\Gamma}}_{2},{{\Gamma}}_{3}$, ${{\Gamma}}_{4}$
$Pm11$ ${{\Gamma}}_{5}$, ${{\Gamma}}_{6}$ ${\Gamma}_{5}$
$Pb11$ ${{\Gamma}}_{1}$, ${{\Gamma}}_{2},{{\Gamma}}_{3}$, ${{\Gamma}}_{4}$ ${{\Gamma}}_{1}$, ${{\Gamma}}_{2},{{\Gamma}}_{3}$, ${{\Gamma}}_{4}$ ${{\Gamma}}_{1}$, ${{\Gamma}}_{2}$ ${{\Gamma}}_{1}$, ${{\Gamma}}_{2},{{\Gamma}}_{3}$, ${{\Gamma}}_{4}$ ${{\Gamma}}_{1}$, ${{\Gamma}}_{2}$
Table 2: Spinless TR-breaking Cases for Symmetry-enforced Moiré Topology. List of combinations of atomic point group $G_{0}^{A}$, atomic $\Gamma$ irrep, and moiré plane gorup $G_{0}$ that lead to symmetry-enforced nontrivial moiré topology, without TR symmetry and without SOC. The meaning of the table entries is the same as that in Tab. 1.
IV Symmetry-Enforced Moiré Topology in 2D Plane Group
The symmetry-enforced moiré topology is not limited to the symmetry data discussed in Sec. II. In this section, we will show that there are many other cases where nontrivial moiré topology can be enforced by the atomic symmetry data and the moiré symmetry group. We still focus on the $\Gamma$-valley electron gas (both with and without SOC and TR) in the weak moiré region (i.e., the moiré gap is weaker than atomic band splitting if present).
As discussed in Sec. II, atomic symmetry data includes the atomic symmetry group $G^{A}$, and its irrep $\Lambda^{A}_{\Gamma}$ furnished by the considered states at $\Gamma$, and we use $G$ to label the moiré symmetry group. Here $G^{A}$ is simply the point group at $\Gamma$, $G^{A}_{0}$, if there is no TR symmetry, but also includes the TR symmetry, denoted as $\mathcal{T}$, if it appears. Similarly, $G$ is simply the moiré plane group, $G_{0}$, if there is no TR symmetry, but also includes the TR symmetry if it appears. We focus on the case that the moiré potential does not cause extra TR breaking. For the specific case in Sec. II, we have $G^{A}=G^{A}_{0}\rtimes\mathcal{T}$ with $G^{A}_{0}=C_{6v}$, the irrep $\Lambda^{A}_{\Gamma}=\bar{\Gamma}_{7}$, and $G=G_{0}\rtimes\mathcal{T}$ with $G_{0}=P6mm$. Although we have the atomic symmetry group to be included by the moiré symmetry group in Sec. II, we do not assume such a relation in the general study in this section. The detailed steps for this general search are illustrated in SM Sec.A [SM].
We went through all possible combinations of $G^{A}$, the irrep $\Lambda^{A}_{\Gamma}$, $G$, and the presence/absence of SOC, as elaborated in SM Sec.B [SM]. The results are summarized in Tab. 1 for the spinful case with TR and in Tab. 2 for the spinless case without TR, where spinful (spinless) simply refers to the case with (without) SOC as indicated by the irrep notation with (without) a bar [aroyo2011crystallography, aroyo2006bilbao, aroyo2006bilbao1]. (The results of spinless cases with TR and spinful cases without TR are summarized in Tab. SI(c) and SIV(b) in the SM Sec.B [SM] respectively. ) We find 58 combinations that lead to symmetry-enforced moiré topology in the spinful case, and 139 in the spinless case. Out of the 58 spinful combinations (139 spinless combinations), we find that there are 11 (5) combinations that allow for isolated topological bands at low-energy, while all other combinations only have topological semi-metal phases in the weak moiré potential limit. The symmetry-enforced gapped moiré topology has the following scenarios: (i) for the TR-preserving case, isolated topological moiré bands may carry nonzero $\mathds{Z}_{2}$ index in addition to other possibilities such as fragile topology; (ii) for the TR-breaking case, the topological moiré bands always have nonzero Chern number.
From our results, we can immediately provide several guiding principles for finding moiré topological insulators in $\Gamma$-valley moiré systems. First, symmetry-enforced moiré gapped topology only happens for doublets at $\Gamma$ for both the TR-preserving and TR-breaking cases. Second, almost all symmetry-enforced moiré topological insulators happen on a triangular/hexagonal lattice, with only three combinations as exceptions. Therefore, to look for moiré topological insulators in $\Gamma$-valley moiré systems, it is favorable to look for low-energy 2D irreps at $\Gamma$ valley subject to triangular/hexagonal moiré potential. Finally, symmetry-enforced moiré topological insulators only happen when SOC and TR symmetry are simultaneously present or simultaneously absent, meaning that the simultaneous presence and absence of TR symmetry and SOC is also favorable for moiré topological insulators in $\Gamma$-valley moiré systems.
On the other hand, symmetry-enforced moiré topological semimetals are much easier to be realized than symmetry-enforced moiré topological insulators. There are no particular constraints on the lattice type, or the simultaneous presence/absence of TR symmetry and SOC. Nevertheless, doublets at $\Gamma$ are still required to realize symmetry-enforced moiré topological semimetal, unless the moiré plane group is non-symmorphic [michel1999connectivity, michel2001elementary, watanabe2015filling, watanabe2016filling, vergniory2017graph].
V Material candidates
We now discuss the material candidates. As our framework only needs information of the monolayers, we can screen through the existing non-moiré 2D material databases [haastrup2018computational, vergniory2019complete, jiang20242d] for suitable monolayer 2D materials for our proposal.
We first pick out monolayer spin-orbit coupled materials in which (i) VBM or CBM is located at $\Gamma$ and (ii) there are no symmetries that will force the atomic band splitting to be zero along at least one line in the BZ. Particularly, we focus on atomic plane groups P31m, P3m1 and P3, in which 92 monolayer materials are identified with their atomic symmetry data listed in Tab. 1. In particular, for 71 of them, the required moiré symmetry group is P3 according to Tab. 1, which can be naturally realized by a small angle twisted homobilayer structure.
We show two examples, BiBrTe and BiFTe, in Fig. 4. For these two examples, the atomic symmetry group at $\Gamma$ is $G_{0}^{A}=C_{3v}$ and the irreps of VBM are $\bar{\Gamma}_{4}\bar{\Gamma}_{5}$, which belong to Tab. 1. The required moiré symmetry group P3 naturally arise in the twist homobilayer of either of the two examples (See more discussion in Sec.D in SM [SM].
Figure 4: (a) Band dispersion of BiBrTe. (b) Band dispersion of BiFTe. The irreps of VBM for both materials are $\bar{\Gamma}_{4}\bar{\Gamma}_{5}$. The inset figures show a zoom-in of the regions inside the dashed boxes.
VI Conclusion and Discussion
We have proposed the concept of symmetry-enforced moiré topology, and list all combinations of atomic symmetry data and moiré symmetry group to realize it for $\Gamma$-valley electron gas subject to a weak moiré potential. 92 candidate monolayer materials are identified based on a selective search.
We note that the TR-breaking cases with the spinless models in Sec.C2 of SM [SM] can also be applied to topological moiré magnon systems [li2020moire, wang2020stacking, ganguli2023visualization]. Furthermore, the generation of the moiré potential is not limited to twisting two homobilayers—it can also be done (i) by forming hetero-structure with insulating materials such as boron nitride or transition metal dichalcogenides (TMD) [yasuda2021stacking, woods2021charge, zhao2021universal, kim2024electrostatic] and (ii) by fabricating patterned hole arrays in dielectric substrate materials [forsythe2018band, barcons2022engineering]. We note that although the homobilayer structure contains two copies of the $\Gamma$-valley modes from each layer, the low-energy physics can only involve either the bonding combination or the anti-bonding combination between these two copies, owing to the strong uniform interlayer coupling [angeli2021gamma, zhang2021electronic, jia2024moire, liu2025ideal]. As a result, at low energies, homobilayer twist structure effectively has only one set of the $\Gamma$-valley modes. This is exactly the case in twisted homobilayer TMD at $\Gamma$ valley [angeli2021gamma, zhang2021electronic, jia2024moire]. While our current study focuses on layered materials with low-energy $\Gamma$-valley modes, the methodology should be generalizable to other valleys, such as $K$ or $M$ valleys [cualuguaru2025moire], which we leave for the future work.
VII Acknowledgment
We acknowledge the helpful discussion with Jennifer Cano and Lei Chen. J. Y.’s work is supported by startup funds at University of Florida. Y.L. and C.L. acknowledge the support from the Penn State Materials Research Science and Engineering Center for Nanoscale Science under National Science Foundation award DMR-2011839. This work was performed in part at Aspen Center for Physics, which is supported by National Science Foundation grant PHY-2210452.